OpenStudy (anonymous):

In a game that lasts 48 Minutes, exactly 6 players from each team are on the field at all times. Throughout the game, players are substituted so that 8 players on a team each play an equal amount of time. How many minutes is each of the 8 players on the field during the game?

OpenStudy (anonymous):

each player therefore plays \[\frac{6}{8}=\frac{3}{4}\] of the time, so compute \[\frac{3}{4}\times 48\]

OpenStudy (anonymous):

so 6?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

you have to multiply \[\frac{3}{4}\times 48\]